OpenStudy (anonymous):

\[Substitute~~~~~\int\limits {1 \over (3x+4)^2} dx ---->3x+4=u\]

OpenStudy (mimi_x3):

Where are you having problems? u = 3x +4 => du/3 It's pretty easy..

OpenStudy (anonymous):

Can you show the steps of how to integrate it?\[\int\limits {1 \over u^2}dx\] becomes....

OpenStudy (mimi_x3):

Write it as \[\int\limits u^{-2} dx\]

OpenStudy (mimi_x3):

are you able to integrate it?

OpenStudy (anonymous):

Yes\[-->{ 1 \over u}\]

OpenStudy (anonymous):

But why is the answer 1/-3(3+4) and not 1/3x+4?

OpenStudy (anonymous):

sorry, it's -1/u

OpenStudy (anonymous):

Why?

OpenStudy (mimi_x3):

\[\int\limits\frac{1}{(u)^{2}} *\frac{du}{3} => \frac{1}{3} \int\limits\frac{1}{u^{2}} \]

OpenStudy (mimi_x3):

You move the 1/3 outside..

OpenStudy (mimi_x3):

then integrate it..

OpenStudy (anonymous):

\[\int\limits \sin({1\over3}\pi-{1\over2}x)dx~~~~~~~Why~would~2~go~on~the~outside~instead~of~1/2?\]

OpenStudy (mimi_x3):

Sorry, I don't know, lol. i forgot this... Someone else should help, lol

OpenStudy (anonymous):

Thanks for trying to help anyway! ;D